OpenStudy (softballgirl372015):

Please help! How do you solve for x? 4sin(x)=cos(x)-2

OpenStudy (anonymous):

let y=sin(x), sqrt(1-y^2)=cos(x). solve for y, then x=arcsin(y)+2*pi*C

OpenStudy (softballgirl372015):

How did you come to that?

OpenStudy (anonymous):

since it is difficult to solve for x, solve for sin(x) instead. then x=arcsin(sin(x))+2*pi*C

OpenStudy (softballgirl372015):

I have not learned arcsin yet. Is there a way to solve this problem without arcsin and maybe patharoyean identities?

OpenStudy (anonymous):

using the fact that sin^2+cos^2=1, you can reduce the equation down to 4 sin=sqrt(1-sin^2)-2 (4 sin+2)=sqrt(1-sin^2) 16 sin^2+16 sin+4=1-sin^2 17 sin^2+16sin+3=0 sin x=(sqrt(13)-8)/17 beyond this point, arcsin is necessary to find x

OpenStudy (softballgirl372015):

In the first line, How did you get the expression under the radical since it is cos, not \[\cos^{2}\]? I attached what i've done, but when I use the inverse cos button on my calc, it says doamin error and I dont know why it wont work.

OpenStudy (anonymous):

sin^2+cos^2=1 cos^2=1-sin^2 cos=sqrt(1-sin^2)

OpenStudy (anonymous):

(cos(x)-2)^2=cos^2+4-4cos(x)

OpenStudy (anonymous):

i think you missed a term

OpenStudy (softballgirl372015):

okay. I am going to try and redo it with what you suggested. What do you think the answer is though?

OpenStudy (anonymous):

in radians or in degrees? in radians, -0.261+2*pi*n where n is an integer in degrees, -14.981+360*n, where n is an integer arcsin((sqrt(13)-8)/17)